# Exceptional supercuspidal representations in small residue characteristic

### v1

@@ 1 Introduction
We now describe our main result. In Section 3, we review the construction of supercuspidal representations due to [RY14] and verify that they are epipelagic. In Definition 3.12 we define a weaker notion of stability than the notion from Geometric Invariant Theory, that of $\mathbb{F}_{q}$-stability. In short, this asks for the Hilbert–Mumford criterion ([Mum77]) to hold for the action of any one parameter family $\mathbb{G}_{m}\to\mathbb{G}_{x}$, defined over $\mathbb{F}_{q}$, on $V(x)^{\vee}$. We prove the following:
Theorem 1.1 (Theorem 3.17).
@@
In Section 4, we give examples of supercuspidal representations using Theorem 3.17. These examples have the following features:
@@
  • For $G$ of type $B_{n}$, $q=2$ and $x$ the barycentre of an alcove, we produce irreducible supercuspidal representations of $G(k)$, of which at least one does not arise from the construction of [RY14] by depth considerations.
@@ 2.2 The abelianisation of $G(k)_{x,+:1}$
This is the pro-unipotent radical of an Iwahori subgroup of $\mathrm{SL}_{2}$ or $\mathrm{PGL}_{2}$. By Lemma 2.1, the derived subgroup of this contains $\langle U_{\alpha-\lceil\alpha(x)\rceil+2},U_{\lceil\alpha(x)\rceil-\alpha+1},\alpha^{\vee}(1+\varpi\mathcal{O}_{k})\rangle.$
@@
This is the pro-unipotent radical of a maximal parahoric subgroup of $\mathrm{SL}_{2}$ or $\mathrm{PGL}_{2}$ (so is conjugate to the kernel of reduction modulo $\varpi$ of $\mathrm{SL}_{2}(\mathcal{O}_{k})$ or $\mathrm{PGL}_{2}(\mathcal{O}_{k})$). By the same calculation as in Lemma 2.1 (replacing $b$ by $\varpi b$), the derived subgroup of this contains $\langle U_{\alpha-\lceil\alpha(x)\rceil+2},U_{\lceil\alpha(x)\rceil-\alpha+2},\alpha^{\vee}(1+\varpi^{2}\mathcal{O}_{k})\rangle.$
@@
To determine the abelianisation of $G(k)_{x,+:1}$ we must describe the derived subgroup. Note that if $0<\psi(x),\phi(x)<1$ and $\psi+\phi$ is constant (necessarily 1), then $U_{\psi},U_{\phi}]\subset\langle U_{\psi+1},U_{\phi+1},T_{1}\rangle\leq G(k)_{x,1}$ by the calculation of Lemma 2.1. Chevalley’s commutator formula for affine roots [Ste16, Lemma 15] gives, for affine roots $\psi,\phi$ with $\psi+\phi$ non-constant and $0<\psi(x),\phi(x)<1$,
@@
As the Killing form is positive definite, by Lemma 2.4 we must have that $\dot{\psi}$ pairs positively with $\dot{\theta}$ for some $\theta\in\bigcup\limits_{w\in W(x)}w(\Pi(x))\subset\Delta(x)$, with respect to $(,)$. Since $0<\theta(x)<\psi(x)\leq 1$, we know that $\dot{\psi}\neq\dot{\theta}$. So $\psi-\theta\in\Psi$, and $0<(\psi-\theta)(x)\leq 1$. ∎
Notation 2.8.
@@
Suppose $1\geq\psi(x)\geq\delta(x)m>\delta(x)$. Then $\psi\not\in\Delta(x)$. By Corollary 2.7 there exists $\theta\in\Delta(x)$ such that $\psi-\theta\in\Psi_{x}^{+}$. Applying the commutator relations to $\theta$ and $\psi-\theta$, we see that
@@
    for all $r,s\in\mathcal{O}_{k}$. By Lemma 2.3, $C_{\dot{\theta},\dot{\psi}-\dot{\theta},1,1}=\pm 1$, and we deduce that $\overline{U}_{\psi}\leq(G(k)_{x,+:1})_{\mathrm{der}}$.
@@
Suppose $\dot{\psi}$ is short. Then by Corollary 2.7, there exists $\theta\in\Delta(x)$ such that $\langle\dot{\psi},\dot{\theta}^{\vee}\rangle>0$, which is necessarily 1 as $\dot{\psi}$ is short. If $\langle\dot{\theta},\dot{\psi}^{\vee}\rangle=1$, then $\overline{U}_{\psi}\leq(G(k)_{x,+:1})_{\mathrm{der}}$ by Lemma 2.11. Otherwise, $2\dot{\psi}-\dot{\theta}\in\Phi$. Since $\psi\not\in\Delta(x)$, we have $(2\psi-\theta)(x)>\psi(x)\geq m\delta(x)$. If $2\psi-\theta\not\in S(x)$, then the induction hypothesis implies that $\overline{U}_{2\psi-\theta}\leq(G(k)_{x,+:1})_{\mathrm{der}}$, and this forces $\overline{U}_{\psi}\leq(G(k)_{x,+:1})_{\mathrm{der}}$. Otherwise, $2\psi-\theta$ is of the form $\psi_{1}+\psi_{2}$ or $2\psi_{1}+\psi_{2}$ for $\psi_{1},\psi_{2}\in\Delta(x)$. In the first case, $\psi-\theta\in\Psi$ takes value $\delta(x)/2$ at $x$, contradicting the minimality of $\delta(x)$. In the second case, the equality $(\psi-\theta)(x)=\delta(x)$ means that $\psi-\theta\in\Delta(x)$ by definition. But $\psi=\theta+(\psi-\theta)$ contradicts our assumption that $\psi\not\in S(x)$.
@@
Suppose now $\dot{\psi}$ is long. Then by Corollary 2.7, there exists $\theta\in\Delta$ such that $\langle\dot{\psi},\dot{\theta}^{\vee}\rangle\in\{1,2\}$. If the value is 1, the same argument as above holds to show that $\overline{U}_{\psi}\leq(G(k)_{x,+:1})_{\mathrm{der}}$. If the value is 2 then the commutator formula gives
$u_{\psi-2\theta}(r),u_{\theta}(s)]=u_{\psi-\theta}((-r)s)\cdot u_{\psi}((-r)s^{2})\mod G(k)_{x,1}$ (1)
@@
for all $r,s\in\mathcal{O}_{k}$. Consider first the case that $\psi-\theta\in S(x)$. If $\psi-\theta\in\Delta(x)$ then $\psi=\theta+(\psi-\theta)\in S(x)$. If $\psi-\theta=\psi_{1}+\psi_{2}$ for some $\psi_{1},\psi_{2}\in\Delta(x)$, then $\psi-2\theta\in\Psi$ takes value $\delta(x)$ at $x$, so is in $\Delta(x)$, and so $\psi=2\theta+(\psi-2\theta)\in S(x)$. The case that $\psi-\theta=2\psi_{1}+\psi_{2}$ for some $\psi_{1},\psi_{2}\in\Delta(x)$ does not occur as the gradient of $\psi-\theta$ is a short root, and the gradient of $2\psi_{1}+\psi_{2}$ is a long root. In any case, we have a contradiction to $\psi\not\in S(x)$. Finally, consider the case that $\psi-\theta\not\in S(x)$. The gradient $\dot{\psi}-\dot{\theta}$ is a short root, and again by Corollary 2.7 there exists $\phi\in\Delta(x)$ such that $\langle\dot{\psi}-\dot{\theta},\dot{\phi}^{\vee}\rangle=1$. If $\dot{\phi}$ is a short root, then Lemma 2.11 implies that $\overline{U}_{\psi-\theta}\leq(G(k)_{x,+:1})_{\mathrm{der}}$. Then $\overline{U}_{\psi}\leq(G(k)_{x,+:1})_{\mathrm{der}}$ by (1). Otherwise, $2\dot{\psi}-2\dot{\theta}-\dot{\phi}\in\Phi$. We have $(2\psi-2\theta-\phi)(x)=(2m-3)\delta(x)$. If $m>3$ then $\overline{U}_{2\psi-2\theta-\phi}\leq(G(k)_{x,+:1})_{\mathrm{der}}$ by the induction hypothesis, and then Lemma 2.11 forces $\overline{U}_{\psi-\theta}\leq(G(k)_{x,+:1})_{\mathrm{der}}$, and again we have $\overline{U}_{\psi}\leq(G(k)_{x,+:1})_{\mathrm{der}}$. The case $m=2$ implies $\psi-\theta\in\Delta(x)$, contradicting $\psi-\theta\not\in S(x)$. The case $m=3$ implies $\psi-2\theta\in\Delta(x)$, and then $\psi=2\theta+(\psi-2\theta)$ contradicts $\psi\not\in S(x)$. ∎
@@
$0=\sum\limits_{i,j>0:i\phi+j\theta\in\Psi_{x,<1}^{+}}\nu_{i\phi+j\theta}(C_{\dot{\phi},\dot{\theta},i,j}(-r)^{i}s^{j})$ (2)
in $\mathcal{V}(x)$, for all $r,s\in\mathcal{O}_{k}$. As $\nu_{\psi}$ is trivial when $\psi$ does not appear in $\mathcal{S}(x)$ (as a singleton or a tuple), it suffices to check that (2) holds whenever $\phi,\theta\in\Psi_{x,<1}^{+}$ are such that there exists $i,j>0$ such that $i\phi+j\theta$ appears in $\mathcal{S}(x)$. For any $\psi$ appearing in $\mathcal{S}(x)$ we have $\psi(x)\leq 3\delta(x)$. So we must have $\phi,\theta\in\Delta(x),\Delta_{2}(x)$. Our tedious construction of $\mathcal{S}(x)$ was designed exactly so that (2) holds for all $\phi,\theta\in\Delta(x),\Delta_{2}(x)$. As $\nu_{\psi}$ is surjective when $\psi$ appears in $\mathcal{S}(x)$, the homomorphism $\nu$ is surjective. Any homomorphism $\nu^{\prime}$ from $G(k)_{x,+}$ to an abelian group satisfies the analogue of (2), so that by construction $\nu^{\prime}$ must factor through $\nu$. This is the universal property defining $\mathcal{V}(x)$ as the abelianisation of $G(k)_{x,+:1}$. ∎
Example 2.16.
@@ 3.2 Epipelagic representations
where the last line follows from Frobenius reciprocity. The above sum being nonzero is equivalent to there existing $g$ such that $\prescript{g}{}{\chi_{\lambda}}$ is trivial on $\prescript{g}{}{G(k)_{x,+}}\cap G(k)_{y,r+}$. By the affine Bruhat decomposition, Proposition 3.4, we may write $g=anb$ for $a\in G(k)_{y,0}$, $b\in G(k)_{x,0}$ and $n\in N(k)$. Then $\prescript{b}{}{\chi_{\lambda}}=\chi_{b\cdot\lambda}$ is trivial on
@@
By the elementary lemma following this proof, there then exist $a_{\psi}\geq 0$, not all zero, such that
@@
Lemma 3.11.
Let $V$ be a finite-dimensional vector space over $\mathbb{R}$ and suppose $f_{1},\dots,f_{n}$ are linear functionals on $V$ with the property that for all nonzero $v\in V$, there exists $i$ such that $f_{i}(v)>0$. Then there exist real numbers $a_{i}\geq 0$, not all zero, such that $\sum_{i}a_{i}f_{i}=0$.
Proof.
[Dav54, Theorem 3.1]. ∎
@@ 3.3 Towards higher depth
Lemma 3.13.
@@
Proposition 3.14.
Let $\lambda\in V(x)^{\vee}$ be an $\mathbb{F}_{q}$-stable functional. Fix a positive real number $r$. There exists finitely many $g\in G(k)_{x,r}\backslash G(k)/G(k)_{x,+}$ such that $\lambda$ vanishes on $G(k)_{x,+}\cap g^{-1}G(k)_{x,r}g$.
@@
By Proposition 3.4 a set of $G(k)_{x,0}$ double coset representatives of $G(k)$ is contained in a set of lifts of elements of $W(G)$ to $N(k)$. Write $g=anb$ where $a,b\in G(k)_{x,0}$ and $n\in N(k)$ has image $w\in W(G)$. Then
@@
Observe that $G(k)_{x,+}\cap G(k)_{y,r}$ contains $U_{\psi}$ for all $\psi\in\Psi$ such that $\psi(x)>0$ and $\psi(y)\geq r$. If $\psi\in\mathrm{supp}(b\cdot\lambda)$ then we automatically have $\psi(x)>0$. So, for such $\psi$, if $\langle\dot{\psi},\gamma\rangle\geq r$, we certainly have $\psi(y)\geq r$. It suffices to show that there exist finitely many $w\in W(G)$ such that for $\gamma=w^{-1}(x)-x$ we have $\langle\dot{\psi},\gamma\rangle\leq r$ for all $\psi\in\mathrm{supp}(b\cdot\lambda)$. This follows from Lemma 3.13 and the fact that under the action of $W(G)$ on $X_{*}(T)\otimes\mathbb{R}$, the stabiliser of any point is finite and the orbit of any point is discrete. ∎
Corollary 3.15.
@@
As $\chi_{\lambda}$ is a character, and $\chi$ is nontrivial on $\mathbb{F}_{q}$, $\mathrm{Hom}_{G(k)_{x,+}\cap G(k)_{x,r}^{g}}(\mathbbm{1}_{G(k)_{x,+}\cap G(k)_{x,r}^{g}},\chi_{\lambda}\mid_{G(k)_{x,+}\cap G(k)_{x,r}^{g}})$ is nonzero (and 1-dimensional) if and only if $\lambda$ vanishes on $G(k)_{x,+}\cap G(k)_{x,r}^{g}$. The result follows from Proposition 3.14. ∎
Theorem 3.17.
@@
This follows immediately from Corollary 3.16 and [Bus90, Theorem 1]. ∎
@@
Suppose that $\psi\in\mathrm{supp}(b\cdot\lambda)=\mathrm{supp}(\chi_{\lambda}^{b^{-1}})$ and $w(\psi)\in\Psi_{x}^{+}$ but $w(\psi)\not\in\mathrm{supp}(a\cdot\lambda)=\mathrm{supp}(\chi_{\lambda}^{a^{-1}})$. If $a^{-1}nb$ intertwines $\chi_{\lambda}$, then $\chi_{\lambda}=\chi_{\lambda}^{a^{-1}nb}$ on $G(k)_{x,+}\cap G(k)_{x,+}^{a^{-1}nb}$. So $\chi_{\lambda}^{b^{-1}}=\chi_{\lambda}^{a^{-1}n}$ on $G(k)_{x,+}^{b^{-1}}\cap G(k)_{x,+}^{a^{-1}n}=G(k)_{x,+}\cap G(k)_{x,+}^{n}$. This equality follows from the fact that $G(k)_{x,+}$ is normal in $G(k)_{x,0}$. Since $U_{\psi}=U_{w(\psi)}^{n}$, and $w(\psi)\in\Psi_{x}^{+}$, we have $U_{\psi}\leq G(k)_{x,+}\cap G(k)_{x,+}^{n}$. By assumption, $\chi_{\lambda}^{a^{-1}}$ is trivial on $U_{w(\psi)}$, so $\chi_{\lambda}^{b^{-1}}=\chi_{\lambda}^{a^{-1}n}$ is trivial on $U_{\psi}=U_{w(\psi)}^{n}$, a contradiction to $\psi\in\mathrm{supp}(\chi^{b^{-1}})$. ∎
Lemma 3.19.
@@
Suppose such a $y$ existed. Then for all $\psi\in\mathrm{supp}(b\cdot\lambda)$, we have that $w(\psi)\not\in\Psi_{x}^{+}$ (and so $w(\psi)$ is non-positive on $\mathcal{F}$) or $\psi(w^{-1}y)=w(\psi)(y)=0$. By definition, $b\cdot\lambda$ is still $\mathbb{F}_{q}$-stable. By Lemma 3.11, there exist real numbers $a_{\psi}\geq 0$ such that

### v2

@@ 1 Introduction
We now describe our main result. In Section 3, we review the construction of supercuspidal representations due to [RY14] and verify that they are epipelagic. In Definition 3.3.1 we define a weaker notion of stability than the notion from Geometric Invariant Theory, that of $\mathbb{F}_{q}$-stability. In short, this asks for the Hilbert–Mumford criterion ([Mum77]) to hold for the action of any one parameter family $\mathbb{G}_{m}\to\mathbb{G}_{x}$, defined over $\mathbb{F}_{q}$, on $V(x)^{\vee}$. We prove the following:
Theorem 1.0.1 (Theorem 3.3.7).
@@
In Section 4, we give examples of supercuspidal representations using Theorem 3.3.7. These examples have the following features:
@@
  • For $G$ of type $B_{n}$, $q=2$ and $x$ the barycentre of an alcove, we produce irreducible supercuspidal representations of $G(k)$, of which at least one does not arise from the construction of [RY14] by depth considerations. However, we show how these may be obtained from the setting of Reeder–Yu using our weaker notion of $\mathbb{F}_{q}$-stability.
@@ 2.2 The abelianisation of $G(k)_{x,+:1}$
This is the pro-unipotent radical of an Iwahori subgroup of $\mathrm{SL}_{2}$ or $\mathrm{PGL}_{2}$. By Lemma 2.2.1, the derived subgroup of this contains $\langle U_{\alpha-\lceil\alpha(x)\rceil+2},U_{\lceil\alpha(x)\rceil-\alpha+1},\alpha^{\vee}(1+\varpi\mathcal{O}_{k})\rangle.$
@@
This is the pro-unipotent radical of a maximal parahoric subgroup of $\mathrm{SL}_{2}$ or $\mathrm{PGL}_{2}$ (so is conjugate to the kernel of reduction modulo $\varpi$ of $\mathrm{SL}_{2}(\mathcal{O}_{k})$ or $\mathrm{PGL}_{2}(\mathcal{O}_{k})$). By the same calculation as in Lemma 2.2.1 (replacing $b$ by $\varpi b$), the derived subgroup of this contains $\langle U_{\alpha-\lceil\alpha(x)\rceil+2},U_{\lceil\alpha(x)\rceil-\alpha+2},\alpha^{\vee}(1+\varpi^{2}\mathcal{O}_{k})\rangle.$
@@
To determine the abelianisation of $G(k)_{x,+:1}$ we must describe the derived subgroup. Note that if $0<\psi(x),\phi(x)<1$ and $\psi+\phi$ is constant (necessarily 1), then $U_{\psi},U_{\phi}]\subset\langle U_{\psi+1},U_{\phi+1},T_{1}\rangle\leq G(k)_{x,1}$ by the calculation of Lemma 2.2.1. Chevalley’s commutator formula for affine roots [Ste16, Lemma 15] gives, for affine roots $\psi,\phi$ with $\psi+\phi$ non-constant and $0<\psi(x),\phi(x)<1$,
@@
As the Killing form is positive definite, by Lemma 2.2.4 we must have that $\dot{\psi}$ pairs positively with $\dot{\theta}$ for some $\theta\in\bigcup\limits_{w\in W(x)}w(\Pi(x))\subset\Delta(x)$, with respect to $(,)$. Since $0<\theta(x)<\psi(x)\leq 1$, we know that $\dot{\psi}\neq\dot{\theta}$. So $\psi-\theta\in\Psi$, and $0<(\psi-\theta)(x)\leq 1$. ∎
Notation 2.2.8.
@@
Suppose $1\geq\psi(x)\geq\delta(x)m>\delta(x)$. Then $\psi\not\in\Delta(x)$. By Corollary 2.2.7 there exists $\theta\in\Delta(x)$ such that $\psi-\theta\in\Psi_{x}^{+}$. Applying the commutator relations to $\theta$ and $\psi-\theta$, we see that
@@
    for all $r,s\in\mathcal{O}_{k}$. By Lemma 2.2.3, $C_{\dot{\theta},\dot{\psi}-\dot{\theta},1,1}=\pm 1$, and we deduce that $\overline{U}_{\psi}\leq(G(k)_{x,+:1})_{\mathrm{der}}$.
@@
Suppose $\dot{\psi}$ is short. Then by Corollary 2.2.7, there exists $\theta\in\Delta(x)$ such that $\langle\dot{\psi},\dot{\theta}^{\vee}\rangle>0$, which is necessarily 1 as $\dot{\psi}$ is short. If $\langle\dot{\theta},\dot{\psi}^{\vee}\rangle=1$, then $\overline{U}_{\psi}\leq(G(k)_{x,+:1})_{\mathrm{der}}$ by Lemma 2.2.11. Otherwise, $2\dot{\psi}-\dot{\theta}\in\Phi$. Since $\psi\not\in\Delta(x)$, we have $(2\psi-\theta)(x)>\psi(x)\geq m\delta(x)$. If $2\psi-\theta\not\in S(x)$, then the induction hypothesis implies that $\overline{U}_{2\psi-\theta}\leq(G(k)_{x,+:1})_{\mathrm{der}}$, and this forces $\overline{U}_{\psi}\leq(G(k)_{x,+:1})_{\mathrm{der}}$. Otherwise, $2\psi-\theta$ is of the form $\psi_{1}+\psi_{2}$ or $2\psi_{1}+\psi_{2}$ for $\psi_{1},\psi_{2}\in\Delta(x)$. In the first case, $\psi-\theta\in\Psi$ takes value $\delta(x)/2$ at $x$, contradicting the minimality of $\delta(x)$. In the second case, the equality $(\psi-\theta)(x)=\delta(x)$ means that $\psi-\theta\in\Delta(x)$ by definition. But $\psi=\theta+(\psi-\theta)$ contradicts our assumption that $\psi\not\in S(x)$.
@@
Suppose now $\dot{\psi}$ is long. Then by Corollary 2.2.7, there exists $\theta\in\Delta$ such that $\langle\dot{\psi},\dot{\theta}^{\vee}\rangle\in\{1,2\}$. If the value is 1, the same argument as above holds to show that $\overline{U}_{\psi}\leq(G(k)_{x,+:1})_{\mathrm{der}}$. If the value is 2 then the commutator formula gives
$u_{\psi-2\theta}(r),u_{\theta}(s)]=u_{\psi-\theta}((-r)s)\cdot u_{\psi}((-r)s^{2})\mod G(k)_{x,1}$ (2.2.13)
@@
for all $r,s\in\mathcal{O}_{k}$. Consider first the case that $\psi-\theta\in S(x)$. If $\psi-\theta\in\Delta(x)$ then $\psi=\theta+(\psi-\theta)\in S(x)$. If $\psi-\theta=\psi_{1}+\psi_{2}$ for some $\psi_{1},\psi_{2}\in\Delta(x)$, then $\psi-2\theta\in\Psi$ takes value $\delta(x)$ at $x$, so is in $\Delta(x)$, and so $\psi=2\theta+(\psi-2\theta)\in S(x)$. The case that $\psi-\theta=2\psi_{1}+\psi_{2}$ for some $\psi_{1},\psi_{2}\in\Delta(x)$ does not occur as the gradient of $\psi-\theta$ is a short root, and the gradient of $2\psi_{1}+\psi_{2}$ is a long root. In any case, we have a contradiction to $\psi\not\in S(x)$. Finally, consider the case that $\psi-\theta\not\in S(x)$. The gradient $\dot{\psi}-\dot{\theta}$ is a short root, and again by Corollary 2.2.7 there exists $\phi\in\Delta(x)$ such that $\langle\dot{\psi}-\dot{\theta},\dot{\phi}^{\vee}\rangle=1$. If $\dot{\phi}$ is a short root, then Lemma 2.2.11 implies that $\overline{U}_{\psi-\theta}\leq(G(k)_{x,+:1})_{\mathrm{der}}$. Then $\overline{U}_{\psi}\leq(G(k)_{x,+:1})_{\mathrm{der}}$ by (2.2.13). Otherwise, $2\dot{\psi}-2\dot{\theta}-\dot{\phi}\in\Phi$. We have $(2\psi-2\theta-\phi)(x)=(2m-3)\delta(x)$. If $m>3$ then $\overline{U}_{2\psi-2\theta-\phi}\leq(G(k)_{x,+:1})_{\mathrm{der}}$ by the induction hypothesis, and then Lemma 2.2.11 forces $\overline{U}_{\psi-\theta}\leq(G(k)_{x,+:1})_{\mathrm{der}}$, and again we have $\overline{U}_{\psi}\leq(G(k)_{x,+:1})_{\mathrm{der}}$. The case $m=2$ implies $\psi-\theta\in\Delta(x)$, contradicting $\psi-\theta\not\in S(x)$. The case $m=3$ implies $\psi-2\theta\in\Delta(x)$, and then $\psi=2\theta+(\psi-2\theta)$ contradicts $\psi\not\in S(x)$. ∎
@@
$0=\sum\limits_{i,j>0:i\phi+j\theta\in\Psi_{x,<1}^{+}}\nu_{i\phi+j\theta}(C_{\dot{\phi},\dot{\theta},i,j}(-r)^{i}s^{j})$ (2.2.17)
in $\mathcal{V}(x)$, for all $r,s\in\mathcal{O}_{k}$. As $\nu_{\psi}$ is trivial when $\psi$ does not appear in $\mathcal{S}(x)$ (as a singleton or a tuple), it suffices to check that (2.2.17) holds whenever $\phi,\theta\in\Psi_{x,<1}^{+}$ are such that there exists $i,j>0$ such that $i\phi+j\theta$ appears in $\mathcal{S}(x)$. For any $\psi$ appearing in $\mathcal{S}(x)$ we have $\psi(x)\leq 3\delta(x)$. So we must have $\phi,\theta\in\Delta(x),\Delta_{2}(x)$. Our tedious construction of $\mathcal{S}(x)$ was designed exactly so that (2.2.17) holds for all $\phi,\theta\in\Delta(x),\Delta_{2}(x)$. As $\nu_{\psi}$ is surjective when $\psi$ appears in $\mathcal{S}(x)$, the homomorphism $\nu$ is surjective. Any homomorphism $\nu^{\prime}$ from $G(k)_{x,+}$ to an abelian group satisfies the analogue of (2.2.17), so that by construction $\nu^{\prime}$ must factor through $\nu$. This is the universal property defining $\mathcal{V}(x)$ as the abelianisation of $G(k)_{x,+:1}$. ∎
Example 2.2.18.
@@ 3.2 Epipelagic representations
where the last line follows from Frobenius reciprocity. The above sum being nonzero is equivalent to there existing $g$ such that $\prescript{g}{}{\chi_{\lambda}}$ is trivial on $\prescript{g}{}{G(k)_{x,+}}\cap G(k)_{y,r+}$. By the affine Bruhat decomposition, Proposition 3.1.4, we may write $g=anb$ for $a\in G(k)_{y,0}$, $b\in G(k)_{x,0}$ and $n\in N(k)$. Then $\prescript{b}{}{\chi_{\lambda}}=\chi_{b\cdot\lambda}$ is trivial on
@@
We have the following elementary lemma:
Lemma 3.2.7.
[Dav54] Let $V$ be a finite-dimensional vector space over $\mathbb{R}$ and suppose $f_{1},\dots,f_{n}$ are linear functionals on $V$ with the property that for all nonzero $v\in V$, there exists $i$ such that $f_{i}(v)>0$. Then there exist real numbers $a_{i}\geq 0$, not all zero, such that $\sum_{i}a_{i}f_{i}=0$.
Proof.
[Dav54, Theorem 3.1]. ∎
As a result, there exist $a_{\psi}\geq 0$, not all zero, such that
@@ 3.3 Towards higher depth
Remark 3.3.2.
At least when $q=2$, this is a strictly weaker condition than stability as defined in the previous subsection. This means that one could use the same method as Reeder–Yu to produce supercuspidal representations coming from vectors in $V_{x}^{\vee}$ that may be $\mathbb{F}_{q}$-stable but not stable. For an example of this, see Section 4.2. The main observation in this example is that the adjoint representation of $\mathrm{SL}_{2}$ over $\overline{\mathbb{F}}_{2}$ has no stable vectors (this is true of any adjoint representation) but, identifying the adjoint representation with the representation on degree 2 homogeneous polynomials in $X,Y$ over $\overline{\mathbb{F}}_{2}$, the action of $\mathrm{SL}_{2}(\mathbb{F}_{2})$ on $X^{2}+XY+Y^{2}$ is trivial. In Section 4.2, for $G$ of type $B_{n}$, we consider a point $y$ in the apartment with $\mathbb{G}_{y}(\mathbb{F}_{2})\cong\mathrm{SL}_{2}(\mathbb{F}_{2})$, from which we produce an $\mathbb{F}_{2}$-stable vector that is not stable.
Lemma 3.3.3.
@@
Proposition 3.3.4.
Let $\lambda\in V(x)^{\vee}$ be an $\mathbb{F}_{q}$-stable functional. Fix a positive real number $r$. There exist finitely many $g\in G(k)_{x,r}\backslash G(k)/G(k)_{x,+}$ such that $\lambda$ vanishes on $G(k)_{x,+}\cap g^{-1}G(k)_{x,r}g$.
@@
By Proposition 3.1.4 a set of $G(k)_{x,0}$ double coset representatives of $G(k)$ is contained in a set of lifts of elements of $W(G)$ to $N(k)$. Write $g=anb$ where $a,b\in G(k)_{x,0}$ and $n\in N(k)$ has image $w\in W(G)$. Then
@@
Observe that $G(k)_{x,+}\cap G(k)_{y,r}$ contains $U_{\psi}$ for all $\psi\in\Psi$ such that $\psi(x)>0$ and $\psi(y)\geq r$. If $\psi\in\mathrm{supp}(b\cdot\lambda)$ then we automatically have $\psi(x)>0$. So, for such $\psi$, if $\langle\dot{\psi},\gamma\rangle\geq r$, we certainly have $\psi(y)\geq r$. It suffices to show that there exist finitely many $w\in W(G)$ such that for $\gamma=w^{-1}(x)-x$ we have $\langle\dot{\psi},\gamma\rangle\leq r$ for all $\psi\in\mathrm{supp}(b\cdot\lambda)$. This follows from Lemma 3.3.3 and the fact that under the action of $W(G)$ on $X_{*}(T)\otimes\mathbb{R}$, the stabiliser of any point is finite and the orbit of any point is discrete. ∎
Corollary 3.3.5.
@@
As $\chi_{\lambda}$ is a character, and $\chi$ is nontrivial on $\mathbb{F}_{q}$, $\mathrm{Hom}_{G(k)_{x,+}\cap G(k)_{x,r}^{g}}(\mathbbm{1}_{G(k)_{x,+}\cap G(k)_{x,r}^{g}},\chi_{\lambda}\mid_{G(k)_{x,+}\cap G(k)_{x,r}^{g}})$ is nonzero (and 1-dimensional) if and only if $\lambda$ vanishes on $G(k)_{x,+}\cap G(k)_{x,r}^{g}$. The result follows from Proposition 3.3.4. ∎
Theorem 3.3.7.
@@
This follows immediately from Corollary 3.3.6 and [Bus90, Theorem 1]. ∎
@@
Suppose that $\psi\in\mathrm{supp}(b\cdot\lambda)=\mathrm{supp}(\chi_{\lambda}^{b^{-1}})$ and $w(\psi)\in\Psi_{x}^{+}$ but $w(\psi)\not\in\mathrm{supp}(a\cdot\lambda)=\mathrm{supp}(\chi_{\lambda}^{a^{-1}})$. If $a^{-1}nb$ intertwines $\chi_{\lambda}$, then $\chi_{\lambda}=\chi_{\lambda}^{a^{-1}nb}$ on $G(k)_{x,+}\cap G(k)_{x,+}^{a^{-1}nb}$. So $\chi_{\lambda}^{b^{-1}}=\chi_{\lambda}^{a^{-1}n}$ on $G(k)_{x,+}^{b^{-1}}\cap G(k)_{x,+}^{a^{-1}n}=G(k)_{x,+}\cap G(k)_{x,+}^{n}$. This equality follows from the fact that $G(k)_{x,+}$ is normal in $G(k)_{x,0}$. Since $U_{\psi}=U_{w(\psi)}^{n}$, and $w(\psi)\in\Psi_{x}^{+}$, we have $U_{\psi}\leq G(k)_{x,+}\cap G(k)_{x,+}^{n}$. By assumption, $\chi_{\lambda}^{a^{-1}}$ is trivial on $U_{w(\psi)}$, so $\chi_{\lambda}^{b^{-1}}=\chi_{\lambda}^{a^{-1}n}$ is trivial on $U_{\psi}=U_{w(\psi)}^{n}$, a contradiction to $\psi\in\mathrm{supp}(\chi_{\lambda}^{b^{-1}})$. ∎
Lemma 3.3.9.
@@
Suppose such a $y$ existed. Then for all $\psi\in\mathrm{supp}(b\cdot\lambda)$, we have that $w(\psi)\not\in\Psi_{x}^{+}$ (and so $w(\psi)$ is non-positive on $\mathcal{F}$) or $\psi(w^{-1}y)=w(\psi)(y)=0$. By definition, $b\cdot\lambda$ is still $\mathbb{F}_{q}$-stable. By Lemma 3.2.7, there exist real numbers $a_{\psi}\geq 0$ such that